OpenStudy (anonymous):

METALS AND FANNING A bag contains 10 white gumballs and 12 orange gumballs. What is the ratio, written in lowest terms, of white gumballs to the total number of gumballs in the bag?

OpenStudy (anonymous):

can someone please help

ganeshie8 (ganeshie8):

As a start, find the total gumballs in the bag

OpenStudy (anonymous):

ok so 22 total

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Yes next take the ratio of `white` gamballs to total balls

OpenStudy (anonymous):

so is that the answer

OpenStudy (anonymous):

10/22

ganeshie8 (ganeshie8):

cancel common factors

ganeshie8 (ganeshie8):

\[\large \dfrac{10}{22} = \dfrac{2\times 5}{2\times 11} =?\]

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

10/22

OpenStudy (anonymous):

10/22=2×5/2×11= 10/22

ganeshie8 (ganeshie8):

cancel 2's

ganeshie8 (ganeshie8):

\[\require{cancel} \large \dfrac{10}{22} = \dfrac{\cancel{2}\times 5}{\cancel{2}\times 11} =\dfrac{5}{11}\]

ganeshie8 (ganeshie8):

5/11 is the final simplified form

OpenStudy (anonymous):

i metaled u